# Graph Residual based Method for Molecular Property Prediction

### v1

@@ 3 Model Description
3.1 GNN Block
A significant amount of work in the field of GNN, [41, 42] focuses on the utilization of node features to find different patterns and relations between different molecule structures, but fails to consider the edge features, which are particularly important when talking about organic molecules, as information such as number of bonds and aromaticity are an important influence on various properties of organic molecules.
Thus, we consider Edge Conditioned Graph Convolution Neural Networks [43], which modifies the normal message passing algorithms of the traditional GNN to include a correction made by using the relevant edge features- aromaticity, multiple bond presence, resonance, etc. Using an Artificial Neural Network, we obtain representations of the edge features in the feature space of the node features, and then add them to the node features during message passing algorithm, ensuring that the resulting node features obtained have information not just from the surrounding neighbouring nodes, but from the edge features connecting those nodes as well.
The Eq.1 describes the convolution operation of the ECC, where $h_{\theta}$ is a multi–layer perceptron which takes in the edge features $(e_{ij})$ and dynamically gives the required edge weight. Here, $x$ represents the node features that are passed to the model, and $\theta$ represents the parameters of the model to be updated.
@@ 3.1 GNN Block
While this solves the inherent problems within implementations of our GNN, we optimize the training with several other methods. It has been observed that many molecules have similar substructures such as aldehyde groups, benzene rings, and other common elements. We believe that since these groups often act in a fairly similar manner, contributing to properties such as solubility, boiling point, etc. in a standard manner across different molecules, a mechanism has been introduced to store these influences in a separate state so that this information can travel through various molecules. Bigger molecules will thus contribute information that can act as a template for smaller molecules, resulting in faster computation efficiency, and a more focused learning. GGNN (Gated Graph Neural Network) [18] tries to store temporal states by induction of a GRU unit in place of each graph node of a GNN during the message passing layer. Since we need a spatial log of the states that are observed by the algorithm, we apply a GRU layer each at the end of the “hops”, i.e. the end of the message passing algorithm, saving the state as needed to preserve node properties across different molecules. The GRU update steps at the $l^{th}$ layer are
[…]
Figure 2: Schematic diagram of the flow of our GNN model.
@@
This approach clearly delineates the flow of information from one layer to the next, emphasizing the sequential dependency between the ECC output and GRU input within the same layer, and the GRU output serving as the ECC input for the subsequent layer.
@@
A pictorial representation of our model is shown in Figure 2 which describes in detail the various parts of our model. The figure shows a combined diagram of our GNN model along with the attached VAE model for classification purposes. As evident from Figure 2 and 9 layers of GNN blocks are added which consist of an ECC layer, Batch Normalization layer, and GRU cell. The number of layers added to the model represents the number of k hops distance from the node, the model aggregates information from. In this case, since we added 3 layers, we are considering each node to aggregate information 3 nodes away. This is especially useful in the case of larger molecules.
@@
However, in our case, too we experiment with deeper models. Experimentation has been done with upto 5 hop structure, but it did not lead to better accuracy of our results, rather the model performed worse. Kipf and Welling [44] showed that increasing the number of hops could decrease the performance due to the inclusion of noise from the additional hop structures. To solve this problem similar to Residual connections in Computer Vision, several methods have been proposed e.g. Highway GCN by Rashimi et al [45], and Jump Knowledge Network by Xu et al [46] .
@@
In this implementation, we use a simple Residual network connection, deriving inspiration from the DeepGCN model of Li et al [47], who used Residual Connections from the Resnet model [48]. A simple version of ResGCN [49] is used in this case where we have a single residual block, which aggregates information from the first hop and second hop which is to be fed in the third hop. Since all molecular graphs are not large, deeper layers are not used. However, a proper accurate comparison of Residual graph networks on molecular structure will be included in a future study.
@@
Additionally, we apply a GraphConv Mechanism [22] described in Eq. 3 after the third hop, with pooling at the top of it. This transforms the node and edge features into a feature space where all the edge features are 1, thus filtering all the remaining information of the edge features into the node features. This is then passed through a pooling layer to obtain a representation of the molecule that can be passed through any subsequent ANN for classification or regression tasks. The implementation of the same is described by
@@
where $W_{1}$ and $W_{2}$ are the parameters of the parameters and $e_{ji}$ is the edge weight which by default we consider it 1. The $x$ here represents the node features computed from the successive ECC blocks.
The ANN defined is a simple Multilayer Perceptron consisting of 3 hidden layers of 128 dimensions each and an output layer. Based on the Regression task the output layer dimension is 1 and for the classification task, the output layer dimension is 256 since the embeddings generated by the model are passed to the VAE as described in Sec. 3.2.
@@ 3.2 Explanation of the GNN algorithm
In the current model, we add a GRU layer after each hop. The addition of the GRU layer along with residual connection is what increases the performance of the model. A logical explanation along with outputs of some explainers like GNNExplainer [50], and PGExplainer [51] has been presented as proofs of the working of the model.
@@
As can be understood from the input, the molecular graphs are dynamic in nature, i.e. their size varies. Due to this dynamic nature of input, the model encounters molecules of different lengths during training and testing. Despite training the model on a random selection of training set, the model may find it difficult to adjust the weights to a varying set of inputs. Thus, to solve this problem and allow better adaptability to the molecules we implement GRU layers after each hop along with residual connection between the first and the third hop.
@@
To understand the detailed working and effectiveness of the GRU and residual connections, GNNExplainer [50] has been used to present an explanation on the behaviour of the model for selected target property. For this we select two datasets, Lipophilicity [52] and Aqueous solubility [52].
@@
Figure 3: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 5-ethylsulfanyl-3-methyl-1-(2-methylpropyl)-6-[[2-(trifluoromethyl)phenyl]methyl] thieno [2,3-d]pyrimidine-2,4-dione, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Lipophilicity.
@@
Figure 4: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 2-[2-[2,4-dimethoxy-5-[(2-methyl-3,4-dihydro-2H-quinolin-1-yl)sulfonyl]anilino]-2-oxoethyl]sulfanylacetic acid, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Lipophilicity.
@@
Figure 5: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 4-methoxy-1H-indazole (COc1cccc2[nH]ncc12), (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Lipophilicity.
@@
Figure 6: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: (4-methylphenyl)urea, (a) Molecular Structure, (b) Node Feature Importance, ,(c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Aqueous Solubility
@@
Figure 7: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 6-(hydroxymethyl)-4-methoxyoxane-2,3,5-triol, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Aqueous Solubility
@@
Figure 8: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 4-methoxy-1H-indazole (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated for the base ECC model, (d) Output generated by our ECC model. In this case the target property considered is Aqueous Solubility.
@@
The Figure 5(c), 5(d) and Figure 6(c), 6(d) show the results of output of the GNNExplainer model. The actual structure of the molecules along with the node importance has been presented in Figure 5(a) and 5(b) which are the top left and top right. The diagrams in 5(c) and 5(d) in the bottom left and bottom right compare the explanation of our model vs the base ECC model as mentioned by [43]. The strength of the node as well as the edge connections are presented by varying the degrees of the black. The atoms as well as the rings are colour-coded to match the molecules.
@@
Lipophilicity and aqueous solubility are properties that describe the degree of solubility of the input substance in non-polar and polar solvents respectively. As the molecules are predominantly organic in nature, the feature importance indicates carbon and nitrogen to be the most important contributing atoms to the determination of these properties. Features such as formal charge and aromaticity are also deemed as important features, as these affect the polarity of the molecule by determining the final charge and polarity of the molecule due to the probabilistic flow of electrons. This, in turn, directly relates to the fact that the more polar a molecule is, the higher the expected solubility in a polar solvent, while high polarity indicates a low solubility in non-polar solvents. An atom (node) being part of a ring is also identified as a feature of importance since that indicates the presence of a cyclic structure which can significantly impact a molecule’s overall shape, electronic distribution, and thus its solubility properties. Cyclic compounds, especially those that are aromatic, tend to have de–localized $\pi$ electrons which contribute to the molecule’s stability and reactivity. This de-localization can affect the molecule’s polarity. Overall, we can determine the interplay of these factors—atom type, formal charge, aromaticity, and structural features such as being part of a ring—determines the lipophilicity and aqueous solubility of organic molecules.
@@
From the explanation graphs of lipophilicity and aqueous solubility, as shown in Figure 5(d) and 8(d), one observation to be noted is that our model focuses intensively on connections that signify a high amount of electron flow. This is visible by the lighter (weaker) connections between bonds of similar atoms, like C-C bonds, and a darker (stronger) connection between bonds of Carbon with other atoms like Oxygen, Nitrogen, Fluorine, etc., which use their inherent electronegativity to draw electrons towards themselves, creating polar bonds. This polarization within the molecule is a key factor in determining its solubility characteristics. Bonds between atoms of different electronegativities (such as C-O, C-N, or C-F bonds) are more polar, making the molecules containing these bonds more likely to interact with polar solvents and less likely to interact with non-polar molecules using dipole-dipole interactions.
@@
In the example of the molecule in Figure 3(a) the performance of our model against the base ECC [43] model, it can be seen that our model accurately captures relations relevant to solubility that are missed by the base ECC models, such as connections within the $CF_{3}$ subgroup, as well as connections within the neighbouring benzene ring. Additionally, our model captures other important polar bonds that affect lipophilicity, which the ECC model fails to capture. Thus, this ablation study provides concrete proof of the validity of our novel ideation. This also explains how having a balance of so-called electron-giving and electron-receiving groups results in a very high lipophilicity. If we observe Figure 6(c) and 6(d), for the molecule (4-methylphenyl) urea we observe that our model captures heterogeneous bonds that are relevant in spreading the electron charge and making the molecule polar, while the base ECC [43] model highlights no such information. Thus, our model accurately captures the reasoning behind the high value of the aqueous solubility of the molecule.
@@ 3.3 Variational Autoencoder (VAE)
Numerous domains, including network system monitoring, proteomic studies, and drug molecule analysis, necessitate the classification of multiple labels based on graph-connected features such as molecule structures and network graphs. Datasets encoded in (SMILES) [39] format, such as Tox21 [52] and Sider [52], encapsulate valuable information pertaining to drug toxicities and other molecule parameters. An initial attempt to establish an enhanced benchmark involved a straightforward model consisting of a single layer of sigmoid neurons appended to the end of a novel GNN, its competitiveness was noteworthy. However, for multilabel prediction tasks, the pertinence of VAEs surpasses conventional models, inspiring a closer exploration of their efficacy.
VAEs operate by encoding input data into a probabilistic latent space and decoding it to reconstruct the input. This reconstruction allows us to align the feature space and the label space, allowing for much more accurate predictions in the multi-label classification scenario.
The utilization of a VAE for capturing the probabilistic latent space presents significant advantages. Firstly, it grants enhanced control over the latent space [15]. Notably, within various AE models, the label-encoder-decoder pathway often outperforms the feature-encoder-decoder pathway. Incorporating the VAE structure within the latent space aids in balancing the complexity of learning and aligning these distinct subspaces. Secondly, achieving smoothness within the latent space is a common objective [16]. VAEs, being probabilistic models, inherently induce local-scale smoothness as the decoder interprets samples rather than specific embeddings. Thirdly, VAE models and their variations exhibit an ability to learn representations characterized by disentangled factors [17]. When both latent spaces, pertaining to features and labels, learn these disentangled factors, it not only facilitates the alignment of these spaces but also contributes significantly to the decoding process, augmenting our results.
@@
While VAEs are often used for generative tasks, the Contrastive Gaussian Mixture Variational Autoencoder (C-GMVAE) [53] is specially designed for multi-label classification (MLC). This model addresses the common challenges faced by vanilla VAEs [54] such as over-regularization, posterior collapse, and label correlation in MLC tasks. The C-GMVAE leverages contrastive learning to strengthen label embedding learning and integrates a Gaussian mixture variational autoencoder (GMVAE) to learn a multimodal latent space synergistically.
The C-GMVAE model comprises two primary components: the Gaussian mixture VAE and the contrastive learning module. The Gaussian mixture VAE adheres to the conventional VAE framework, combining the posterior distribution with a parameter-free isotropic Gaussian prior. During training, it optimizes two essential losses: the KL divergence from the prior to the posterior and the distance between the reconstructed targets and the real targets. This component empowers the model to acquire a multimodal latent space, facilitating probabilistic modeling. A representative diagram of the C-GMVAE model has been presented in Figure 9 to highlight its workflow in the multi-class multi-label classification setting.

### v2

@@ 3 Model Description
3.1 Edge Conditioned Residual Graph Neural Network (ECRGNN)
A significant amount of work in the field of GNNs, 40, 41 focuses on the utilization of node features to find different patterns and relations between different molecule structures but fails to consider the edge features. These are particularly important when talking about organic molecules since information such as the number of bonds and aromaticity influence various properties of organic molecules.
Thus, this paper considers Edge Conditioned Graph Convolution (ECC) Neural Network 42, which modifies the normal message passing algorithms of the traditional GNN to include a correction made by using the relevant edge features- aromaticity, multiple bond presence, resonance, etc. Using an Artificial Neural Network (ANN), representations of the edge features are obtained in the feature space of the node features, and then add them to the node features during message passing algorithm, ensuring that the resulting node features obtained have information not just from the surrounding neighbouring nodes, but from the edge features connecting those nodes as well.
@@ 3.1 Edge Conditioned Residual Graph Neural Network (ECRGNN)
The Eq.1 describes the convolution operation of the ECC, where $h_{\theta}$ is a multi-layer perceptron which takes in the edge features $(e_{ij})$ and dynamically gives the required edge weight. Here, $x$ represents the node features that are passed to the model, and $\theta$ represents the parameters of the model to be updated.
@@
While this solves the inherent problems within implementations of ECRGNN, the training is optimized with several other methods. It has been observed that many molecules have similar substructures, such as aldehyde groups, benzene rings, and other common elements. Since these groups often contribute to properties such as solubility, boiling point, etc., in a standard manner across different molecules, a mechanism has been introduced to store these influences in a separate state so that this information can travel through various molecules. Bigger molecules will thus contribute information that can act as a template for smaller molecules, resulting in faster computation efficiency and more focused learning. Gated Graph Neural Network (GGNN) 17 tries to store temporal states by induction of a GRU unit in place of each graph node of a GNN during the message passing layer. Since there is a need for a spatial log of the states that are observed by the algorithm, a GRU layer is applied each at the end of the “hops”, i.e. the end of the message passing algorithm, saving the state as needed to preserve node properties across different molecules. The GRU update steps at the $l^{th}$ layer are
Update gate:
$z_{i}^{(l)}=\sigma(W_{z}^{(l)}x_{i}^{(l)}+U_{z}^{(l)}h_{i}^{(l-1)}+b_{z}^{(l)}),$ (2a)
Reset gate:
$r_{i}^{(l)}=\sigma(W_{r}^{(l)}x_{i}^{(l)}+U_{r}^{(l)}h_{i}^{(l-1)}+b_{r}^{(l)}),$ (2b)
Candidate hidden state:
$\tilde{h}_{i}^{(l)}=\tanh(W_{h}^{(l)}x_{i}^{(l)}+U_{h}^{(l)}(r_{i}^{(l)}\odot h_{i}^{(l-1)})+b_{h}^{(l)}),$ (2c)
Final hidden state update:
$h_{i}^{(l)}=(1-z_{i}^{(l)})\odot h_{i}^{(l-1)}+z_{i}^{(l)}\odot\tilde{h}_{i}^{(l)}.$ (2d)
@@
This approach clearly delineates the flow of information from one layer to the next, emphasizing the sequential dependency between the ECC output and GRU input within the same layer and the GRU output serving as the ECC input for the subsequent layer.
Figure 2: Schematic diagram of the flow of ECRGNN model.
@@
A pictorial representation of the ECRGNN model is shown in Figure 2, which describes in detail the various parts of the model. The figure shows a combined diagram of the ECRGNN along with the attached VAE model for classification purposes. As evident from Figure 2 and 9, layers of GNN blocks are added, which consist of an ECC layer, Batch Normalization layer, and GRU cell. The number of layers added to the model represents the k-hops distance from the node from which the model aggregates information. In this case, since 3 layers are added, each node is considered to aggregate information 3 nodes away. This is especially useful in the case of larger molecules.
@@
However, in this case, deeper models have also been experimented with. Experimentation has been done with upto 5 hop structure, but it did not lead to better accuracy of the results, rather the model performed worse. Kipf and Welling 43 showed that increasing the number of hops could decrease the performance due to the inclusion of noise from the additional hop structures. Several methods have been proposed to solve this problem similar to Residual connections in Computer Vision, e.g., Highway GCN by Rashimi et al. 44, and Jump Knowledge Network by Xu et al. 45.
@@
In this implementation, a simple Residual network connection is used, deriving inspiration from the DeepGCN model of Li et al. 46, who used Residual Connections from the Resnet model 47. A simple version of ResGCN 48 is used in this case, whereas ECRGNN has a single residual block, aggregating information from the first and second hop, which is to be fed in the third hop. Since all molecular graphs are not large, deeper layers are not used. However, a future study will include a proper accurate comparison of Residual graph networks on molecular structure.
@@
Additionally, a GraphConv Mechanism 21 as described in Eq. 3 is applied after the third hop, with pooling at the top of it. This transforms the node and edge features into a feature space where all the edge features are 1, thus filtering all the remaining information of the edge features into the node features. This is then passed through a pooling layer to obtain a representation of the molecule that can be passed through any subsequent ANN for classification or regression tasks. The implementation of the same is described by
@@
where $W_{1}$ and $W_{2}$ are the parameters and $e_{ji}$ is the edge weight which by default is considered as 1. The $x$ here represents the node features computed from the successive ECC blocks.
The ANN is defined as a simple multilayer perceptron consisting of 3 hidden layers of 128 dimensions each and an output layer. Based on the Regression task, the output layer dimension is 1, and for the classification task, the output layer dimension is 256 since the embeddings generated by the model are passed to the VAE as described in Sec. 3.2.
@@ 3.2 Explanation of the GNN algorithm
In the current model, a GRU layer after each hop is added. The addition of the GRU layer, along with the residual connection, increases the model’s performance. A logical explanation along with outputs of some explainers like GNNExplainer 49 and PGExplainer 50 has been presented as proofs of the working of the model.
@@
The molecular graph inputs are dynamic in nature, i.e. their size varies. Due to this, the model encounters molecules of different lengths during training and testing. Despite training the model on a random selection of training set, the model may find it difficult to adjust the weights to a varying set of inputs. Thus, GRU layers are implemented after each hop, along with a residual connection between the first and the third hop, to solve this problem and allow better adaptability to the molecules.
@@
To understand the detailed working and effectiveness of the GRU and residual connections, GNNExplainer 49 has been used to present an explanation on the behaviour of the model for selected target property on two datasets, Lipophilicity 51 and Aqueous solubility 51.
@@
Figure 3: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 5-ethylsulfanyl-3-methyl-1-(2-methyl propyl)-6-[[2-(trifluoromethyl)phenyl]methyl] thieno [2,3-d]pyrimidine-2,4-dione, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated by the ECC model, (d) Output generated by the ECRGNN model. In this case, the target property considered is Lipophilicity. (The grayscale value of the arrows signifies the strength of the relationship between the two nodes, which directly correlates to the predicted property)
@@
Figure 4: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 2-[2-[2,4-dimethoxy-5-[(2-methyl-3,4-dihydro-2H-quinolin-1-yl)sulfonyl]anilino]-2-oxoethyl]sulfanylacetic acid, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated by the ECC model, (d) Output generated by the ECRGNN model. In this case, the target property considered is Lipophilicity. (The grayscale value of the arrows signifies the strength of the relationship between the two nodes, which directly correlates to the predicted property)
@@
Figure 5: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 4-methoxy-1H-indazole (COc1cccc2[nH]ncc12), (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated by the ECC model, (d) Output generated by the ECRGNN model. In this case, the target property considered is Lipophilicity. (The grayscale value of the arrows signifies the strength of the relationship between the two nodes, which directly correlates to the predicted property)
@@
Figure 6: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: (4-methylphenyl)urea, (a) Molecular Structure, (b) Node Feature Importance, ,(c) Output generated bt the ECC model, (d) Output generated by the ECRGNN model. In this case, the target property considered is Aqueous Solubility. (The grayscale value of the arrows signifies the strength of the relationship between the two nodes, which directly correlates to the predicted property)
@@
Figure 7: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 6-(hydroxymethyl)-4-methoxyoxane-2,3,5-triol, (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated by the ECC model, (d) Output generated by the ECRGNN model. In this case, the target property considered is Aqueous Solubility. (The grayscale value of the arrows signifies the strength of the relationship between the two nodes, which directly correlates to the predicted property)
@@
Figure 8: Comparative Representation of the outputs of GNNExplainer on the target Molecule Name: 4-methoxy-1H-indazole (a) Molecular Structure, (b) Node Feature Importance, (c) Output generated by the ECC model, (d) Output generated by the ECRGNN model. In this case, the target property considered is Aqueous Solubility. (The grayscale value of the arrows signifies the strength of the relationship between the two nodes, which directly correlates to the predicted property)
@@
The Figures 3(c), 4(c), 5(c), 6(c), 7(c), 8(c) and Figures 3(d), 4(d), 5(d), 6(d), 7(d), 8(d) show the results of output of the GNNExplainer model on ECC model and ECRGNN models respectively.42 The actual structure of the molecules, along with the node importance, has been presented in Figure 3(a), 4(a), 5(a), 6(a), 7(a), 8(a) and 3(b), 4(b), 5(b), 6(b), 7(b), 8(b) respectively, which are the top left and top right correspondingly. The strength of the node and the edge connections are presented by varying the degrees of the black. The atoms, as well as the rings, are colour-coded to match the molecules.
@@
Lipophilicity and aqueous solubility are properties that describe the degree of solubility of the input substance in non-polar and polar solvents, respectively. As the molecules are predominantly organic in nature, the feature importance indicates carbon and nitrogen to be the most important contributing atoms to the determination of these properties. Features such as formal charge and aromaticity are also deemed important features, as these affect the molecule’s polarity by determining the molecule’s final charge and polarity due to the probabilistic flow of electrons. This, in turn, directly relates to the fact that the more polar a molecule is, the higher the expected solubility in a polar solvent, while high polarity indicates a low solubility in non-polar solvents. An atom (node) being part of a ring is also identified as a feature of importance since that indicates the presence of a cyclic structure which can significantly impact a molecule’s overall shape, electronic distribution, and thus its solubility properties. Cyclic compounds, especially those that are aromatic, tend to have de–localized $\pi$ electrons, which contribute to the molecule’s stability and reactivity. This de-localization can affect the molecule’s polarity. Overall, the interplay of these factors—atom type, formal charge, aromaticity, and structural features such as being part of a ring—determines organic molecules’ lipophilicity and aqueous solubility.
@@
From the explanation graphs of lipophilicity and aqueous solubility, as shown in Figure 5(d) and 8(d), one observation to be noted is that the ECRGNN model focuses intensively on connections that signify a high amount of electron flow. This is visible by the lighter (weaker) connections between bonds of similar atoms, like C-C bonds, and a darker (stronger) connection between bonds of Carbon with other atoms like Oxygen, Nitrogen, Fluorine, etc., which use their inherent electronegativity to draw electrons towards themselves, creating polar bonds. This polarization within the molecule is a key factor in determining its solubility characteristics. Bonds between atoms of different electronegativities (such as C-O, C-N, or C-F bonds) are more polar, making the molecules containing these bonds more likely to interact with polar solvents and less likely to interact with non-polar molecules using dipole-dipole interactions.
@@
In the example of the molecule in Figure 3(a) the performance of the ECRGNN model against the ECC 42 model, it can be seen that the ECRGNN model accurately captures relations relevant to solubility that are missed by the ECC models, such as connections within the $CF_{3}$ subgroup, as well as connections within the neighbouring benzene ring. Additionally, the ECRGNN model captures other important polar bonds that affect lipophilicity, which the ECC model fails to capture. Thus, this ablation study provides concrete proof of the validity of the novel ideation. This also explains how having a balance of so-called electron-giving and electron-receiving groups results in a very high lipophilicity.
Figure 6(c) and 6(d) for the molecule (4-methylphenyl) urea depicts that the ECRGNN model captures heterogeneous bonds that are relevant in spreading the electron charge and making the molecule polar, while the ECC 42 model highlights no such information. Thus, the ECRGNN model accurately captures the reasoning behind the high value of the aqueous solubility of the molecule.
@@ 3.3 Variational Autoencoder (VAE)
Numerous domains, including network system monitoring, proteomic studies, and drug molecule analysis, necessitate the classification of multiple labels based on graph-connected features such as molecule structures and network graphs. Datasets encoded in (SMILES) 38 format, such as Tox21 51 and Sider 51, encapsulate valuable information pertaining to drug toxicities and other molecule parameters. An initial attempt to establish an enhanced benchmark involved a straightforward model consisting of a single layer of sigmoid neurons appended to the end of a novel GNN, its competitiveness was noteworthy. However, for multilabel prediction tasks, the efficacy of VAEs surpasses conventional models, inspiring a closer exploration of their workings.
VAEs encode input data into a probabilistic latent space and decode it to reconstruct the input. This reconstruction allows us to align the feature and label spaces, allowing for much more accurate predictions in the multi-label classification scenario.
The utilization of a VAE for capturing the probabilistic latent space presents significant advantages. Firstly, it grants enhanced control over the latent space. 14 Notably, within various AE models, the label-encoder-decoder pathway often outperforms the feature-encoder-decoder pathway. Incorporating the VAE structure within the latent space aids in balancing the complexity of learning and aligning these distinct subspaces. Secondly, achieving smoothness within the latent space is a common objective. 15 VAEs, being probabilistic models, inherently induce local-scale smoothness as the decoder interprets samples rather than specific embeddings. Thirdly, VAE models and their variations exhibit an ability to learn representations characterized by disentangled factors. 16 When both latent spaces, pertaining to features and labels, learn these disentangled factors, it facilitates the alignment of these spaces and contributes significantly to the decoding process, augmenting the results.
@@
While VAEs are often used for generative tasks, the Contrastive Gaussian Mixture Variational Autoencoder (C-GMVAE) 52 is specially designed for multi-label classification (MLC). This model addresses the common challenges faced by vanilla VAEs 53 such as over-regularization, posterior collapse, and label correlation in MLC tasks. The C-GMVAE leverages contrastive learning to strengthen label embedding learning and integrates a Gaussian mixture variational autoencoder (GMVAE) to synergistically learn a multimodal latent space.
The C-GMVAE model comprises two primary components: the Gaussian mixture VAE and the contrastive learning module. The Gaussian mixture VAE adheres to the conventional VAE framework, combining the posterior distribution with a parameter-free isotropic Gaussian prior. During training, it optimizes two essential losses: the KL divergence from the prior to the posterior and the distance between the reconstructed targets and the real targets. This component empowers the model to acquire a multimodal latent space, facilitating probabilistic modelling. A representative diagram of the C-GMVAE model has been presented in Figure 9 to highlight its workflow in the multi-class multi-label classification setting.